OpenStudy (anonymous):

e is the midpoint of df, de=2x+4, and ef =3x-1. find de, ef, and df

OpenStudy (ajprincess):

 so 2x+4=3x-1 solve for x. plug in the value of x in 2x+4 to find de and in 3x-1 to find ef. df=de+ef =2x+4+3x-1 =5x+3 plug in the value of x in 5x+3 to find df